OpenStudy (anonymous):

differentiate the following function y=(9x^2)+4(sqrt x)/(3x)

OpenStudy (anonymous):

i got the answer to be (18x+2/(sqrt x)/3

OpenStudy (anonymous):

not sure if thats right

OpenStudy (anonymous):

the derivative of first term is 18x and 4(sqrt x)/(3x) is differentiated by the quotient rule whch is d(u/v) =[ (v* du/dx ) -( u*dv/dx)] / v^2 where u and v are functions of x did u use this?

OpenStudy (anonymous):

oh hold on an easier way may be to do the division first 4(sqrt x)/(3x) = (4/3) x^(-0.5) in which case dy/dx = -2/3 x^ -1.5 = -2 / (3x^(3/2) so final answer = 18x -2 / (3x^(3/2)

OpenStudy (anonymous):

still there rmal?

OpenStudy (anonymous):

yes..but im still trying to understnd the function because my java not working right

OpenStudy (anonymous):

use y=9x^2 +(4x^1/2)/3x =9x^2 +(4x^-1/2)/3 dy/dx=18x-(4/6)x^-3/2 dy/dx=18x-(2/3)x^-3/2 or dy/dx=18x-(2/3x^3/2)

OpenStudy (anonymous):

looks ok then- 2 people with same answer

OpenStudy (anonymous):

yeah lol.....

OpenStudy (anonymous):

my textbook answer shows otherwise

OpenStudy (anonymous):

oh !!! - u sure? what answer does it gave?

OpenStudy (anonymous):

i meant my online course says another but i plug in the solutions the step by step by still wrong...i think it might be the 3x^(3/2)

OpenStudy (anonymous):

yes thats it..thats our answer..lol

OpenStudy (anonymous):

ours is definitely correct - i checked it with some software i have (verify)